# Boost 3D Reconstruction using Diffusion-based Monocular Camera Calibration

### v2

@@ 3.1 Preliminaries on Diffusion Model
For any given input image ${\mathbf{x}}$, the corresponding latent code is generated by the VAE encoder: ${\mathbf{z}}={\mathcal{E}}({\mathbf{x}})$. The forward diffusion process incrementally adds noise to these latents following ${\mathbf{z}}_{t}:=\alpha_{t}{\mathbf{z}}+\sigma_{t}{\bm{\epsilon}}$, where ${\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},{\bm{I}})$, and $\alpha_{t}$ and $\sigma_{t}$ are parameters defined by the noise schedule, with $t\sim p_{t}$ representing the time step within the diffusion schedule. The denoising network, denoted as ${\bm{\epsilon}}_{\theta}$, aims to reverse the diffusion process to recover the noise-free latent code $\hat{{\mathbf{z}}}$ from the final noisy latent code ${\mathbf{z}}_{T}$. This is achieved by predicting the noise component ${\bm{\epsilon}}_{\theta}({\mathbf{z}}_{t},t)$ at each diffusion step. The original image ${\mathbf{x}}$ is then reconstructed from this denoised latent code using the VAE decoder as $\hat{{\mathbf{x}}}={\mathcal{D}}(\hat{{\mathbf{z}}})$. The whole diffusion model is optimized by minimizing the denoising score matching objective, defined as $\mathbb{E}_{{\mathbf{z}},{\bm{\epsilon}},t}\left[\|{\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}({\mathbf{z}}_{t},t)\|_{2}^{2}\right]$. This objective measures the squared Euclidean distance between the actual noise ${\bm{\epsilon}}$ and the predicted noise. By minimizing this objective, the denoising network learns to accurately estimate the noise, thereby effectively reversing the diffusion process and reconstructing the original data distribution.
@@ 3.3 Camera Intrinsic Estimation
$f(\mathbf{z}^{\bm{x}})=p(\hat{\mathbf{z}}^{\bm{c}}_{T})\prod_{t=1}^{T}p_{\bm{\theta}}\left(\hat{\mathbf{z}}^{\bm{c}}_{t-1}|\hat{\mathbf{z}}^{\bm{c}}_{t}\right),\vspace{-2mm}$ (3)
@@ 3.4 Downstream 3D vision tasks
Sparse-View 3D Reconstruction &\&& Pose Estimation.
@@ 4.4 More 3D Vision Tasks
Sparse-view 3D Reconstruction &\&& Pose Estimation. We further demonstrate that our estimated camera intrinsics can be effectively applied to sparse-view 3D reconstruction, especially when photos are taken with varying focal lengths. We evaluate the reconstruction results of Wang et al. [73] on our self-captured images with and without our estimated intrinsics. As shown in Tab. 3, the reconstruction loss, represented by the mean relative distance between corresponding points, was reduced by around 20% on four in-the-wild scenes, confirming the effectiveness of using intrinsic cues for reconstruction. Additionally, as indicated in Tab. 10 of the supplementary, pose estimation performance is also improved when intrinsic cues are used. Quatitative results are shown in Fig. 12 and Fig. 13 of the supplementary, reconstructions without intrinsic cues exhibit notable distortions and misalignments, whereas incorporating intrinsic cues significantly improves accuracy and alignment.

### v3

@@ 3.1 Preliminaries on Diffusion Model
[…]
@@ 3.3 Camera Intrinsic Estimation
$f(\mathbf{z}^{\bm{x}})=p(\hat{\mathbf{z}}^{\bm{c}}_{T})\prod_{t=1}^{T}p_{\bm{\theta}}\left(\hat{\mathbf{z}}^{\bm{c}}_{t-1}|\hat{\mathbf{z}}^{\bm{c}}_{t}\right),\vskip-5.69054pt$ (3)
@@ 3.4 Downstream 3D vision tasks
Sparse-View 3D Reconstruction &\& Pose Estimation.
@@ 4.4 More 3D Vision Tasks
Sparse-view 3D Reconstruction &\& Pose Estimation. We further demonstrate that our estimated camera intrinsics can be effectively applied to sparse-view 3D reconstruction, especially when photos are taken with varying focal lengths. We evaluate the reconstruction results of Wang et al. [73] on our self-captured images with and without our estimated intrinsics. As shown in Tab. 3, the reconstruction loss, represented by the mean relative distance between corresponding points, was reduced by around 20% on four in-the-wild scenes, confirming the effectiveness of using intrinsic cues for reconstruction. Additionally, as indicated in Tab. 10 of the supplementary, pose estimation performance is also improved when intrinsic cues are used. Quatitative results are shown in Fig. 12 and Fig. 13 of the supplementary, reconstructions without intrinsic cues exhibit notable distortions and misalignments, whereas incorporating intrinsic cues significantly improves accuracy and alignment.